# From Bugs to Benchmarks: A Comprehensive Survey of Software Defect Datasets

### v1

@@ 4.2.1. Defect Collection
There are six datasets in the survey that use static analysis to collect defects. To this end, the dataset creators run existing or newly created static analyses on candidate code. As an example of the former, Zheng et al. (2021) create D2A by running the static analyzer Infer (inf, 2025) on pairs of commits, looking for detected issues that are present in one commit but absent in the next. As an example of the latter, ManySStuBs4J (Karampatsis and Sutton, 2020) uses a custom static analysis approach that compares the ASTs of candidate buggy and fixed programs and searches for a set of specific bug-fix patterns. An advantage of using static analysis is that, depending on the specific tools or techniques used, one can achieve better precision than strategies such as keyword matching (i.e., fewer non-bugs classified as bugs). The main disadvantage is that relying on static analysis to separate bugs from non-bugs will limit a dataset’s scope to the abilities of the static analysis. Many static analyzers only detect specific classes of bugs (e.g., null pointer bugs), and even more generalist analyzers like Cppcheck (cpp, 2025) will necessarily fail to catch some bugs (Habib and Pradel, 2018).
@@
There are three datasets in the survey that use machine learning techniques as part of their data collection processes. Wang et al. (2024) input candidate patches into LLMs, asking the models to evaluate whether those patches fix software vulnerabilities. Their dataset also illustrates that extra care must be taken to ensure that LLM output is sound, which they achieve by using LLMs in conjunction with static analysis tools and by including only defects where the two agree. Similarly, Zheng et al. (2021) and Wu et al. (2020) use machine learning models to distinguish between defects and non-defects based on semantic information, such as bug reports or commit messages.
@@ 4.2.2. Validation
Manual Inspection Full (Timperley et al., 2024; Liu et al., 2024b; Zhang et al., 2024a; Cui et al., 2024; Yu et al., 2024; Xu et al., 2024; Waseem et al., 2024; Das et al., 2024; Tian et al., 2024; Silva et al., 2024; Lee et al., 2024; Zhao et al., 2023a; Kim and Hong, 2023; Morovati et al., 2023; Applis and Panichella, 2023; Guan et al., 2023; Xiong et al., 2023; Liang et al., 2022; Kim et al., 2022; Zhang et al., 2022b; Paltenghi and Pradel, 2022; Liu et al., 2022; Wendland et al., 2021; Gyimesi et al., 2021; Yuan et al., 2021; Paul et al., 2021; Du et al., 2021; Zhang et al., 2020; Murphy et al., 2020; Widyasari et al., 2020; Bagherzadeh et al., 2020; Wang et al., 2020; Benton et al., 2019; Radu and Nadi, 2019; Riganelli et al., 2019; Madeiral et al., 2019; Ponta et al., 2019; Xu et al., 2019; Saha et al., 2018; Zhang et al., 2018; Liu et al., 2019; Lin et al., 2015; Just et al., 2014; Ku et al., 2007; Chen et al., 2021b; Gunawi et al., 2014; Ho et al., 2023; Xu et al., 2023; de Santana et al., 2024; Azad et al., 2023; Zhao et al., 2023b; Wang et al., 2022b; Shen et al., 2021; Eghbali and Pradel, 2020; Franco et al., 2017; Wang et al., 2017; Tan et al., 2018) 57
Partial (Wang et al., 2024; Zhong et al., 2023; Avula et al., 2023; Song et al., 2022; Akimova et al., 2021; Kim et al., 2021b; Zheng et al., 2021; Makhshari and Mesbah, 2021; Ferenc et al., 2020; Wu et al., 2020; D’Ambros et al., 2012; Zaman et al., 2012; Wang et al., 2023a, b; Gu et al., 2019) 15
Defect Reproduction (Timperley et al., 2024; Yu et al., 2024; Tian et al., 2024; Jimenez et al., 2024; Pramod et al., 2024; Silva et al., 2024; Lee et al., 2024; Zhao et al., 2023a; Kim and Hong, 2023; Saavedra et al., 2024; Brust et al., 2023; Jiang et al., 2023; Applis and Panichella, 2023; Bhuiyan et al., 2023; Xiong et al., 2023; Zhang et al., 2022b; Bui et al., 2022; Song et al., 2022; Wendland et al., 2021; Gyimesi et al., 2021; Yuan et al., 2021; Widyasari et al., 2020; Benton et al., 2019; Tomassi et al., 2019; Riganelli et al., 2019; Madeiral et al., 2019; Saha et al., 2018; Zhang et al., 2018; Gao et al., 2018; Le Goues et al., 2015; Just et al., 2014; Jalbert et al., 2011; Selakovic and Pradel, 2017; Xu et al., 2023) 34
@@
Dataset validation is about ensuring that the defects included in a dataset are actual defects. This is an important step for making a useful dataset, since a high false positive rate (i.e., many non-bugs classified as bugs) can taint further research that uses a dataset. Even so, not all dataset creators take this step, and there are legitimate reasons for doing so. For instance, a dataset might be so large that validation becomes impractical, or a dataset’s data collection approach might already have a low false positive rate. Table 7 shows common validation methodologies and datasets using them.
@@
Some dataset creators, such as Lee et al. (2024), Wang et al. (2020), and Silva et al. (2024), opt to manually inspect the code and description of every bug in their datasets (i.e., full manual inspection). This allows researchers to construct a dataset with a low false positive rate and to make qualitative judgments about what bugs should be included, without spending the time and resources needed for manual defect collection. Full manual inspection is quite popular among the datasets surveyed, with 57 (57/132) datasets employing it. The largest (Tian et al., 2024) and smallest (Kim and Hong, 2023) datasets that employ this strategy have 4,253 and 18 entries respectively, with a median of 260 entries.
Conversely, some dataset creators choose to only manually inspect a subset of the defects in their dataset (i.e., partial manual inspection). This is less common than full inspection, with only 15 (15/132) datasets employing it. However, it is attractive for particularly large datasets where full inspection is infeasible, such as D2A (Zheng et al., 2021), the largest dataset that uses this approach, which has over 1.3 million entries. The smallest such dataset (Zaman et al., 2012) still contains 191 entries, with the median size being 5,565 entries.
Typically, partial inspection has no effect on the actual construction of the dataset, but is performed to demonstrate the accuracy of an automated methodology (Zheng et al., 2021). However, there is at least one case where partial inspection is used to filter out defects from a dataset. Zhong et al. (2023) automatically classify each bug in their dataset into one of multiple fault categories. They then manually analyze 50 bugs from each category to ascertain that category’s accuracy, and only include bugs from categories with over 90% accuracy in the final dataset.
@@
Among our surveyed datasets, 34 (34/132) involve defect reproduction, and only one dataset attempts to reproduce only a subset of its defects, while the rest aim for full reproduction. This process is often facilitated by sources that provide a rich environment and execution context, such as continuous integration (CI) systems. For example, BugSwarm (Tomassi et al., 2019) leverages CI configurations and historical build logs to automatically reproduce bugs in Docker-based environments, thus making the automation of reproduction feasible.
However, automation has limitations. Many defects are not covered by existing tests, or their reproduction may require complex setup steps beyond test execution (e.g., user interface-related bugs in AndroR2 (Wendland et al., 2021)). Still, automated reproduction enables the construction of much larger datasets than manual validation strategies. The largest dataset in this category (Brust et al., 2023) contains approximately 168,000 entries. The median size, 225 entries, is comparable to that of datasets which use full manual inspection.
@@ 4.2.3. Categorization
Manual Categorization (Timperley et al., 2024; Liu et al., 2024b; Zhang et al., 2024a; Jiang et al., 2024; Cui et al., 2024; Yu et al., 2024; Xu et al., 2024; Waseem et al., 2024; Das et al., 2024; Yu and Wang, 2024; Song et al., 2023; Zhao et al., 2023a; Kim and Hong, 2023; Morovati et al., 2023; Applis and Panichella, 2023; Guan et al., 2023; Xiong et al., 2023; Liang et al., 2022; Kim et al., 2022; Zhang et al., 2022b; Chen et al., 2022; Paltenghi and Pradel, 2022; Liu et al., 2022; Yuan et al., 2021; Paul et al., 2021; Kamienski et al., 2021; Du et al., 2021; Makhshari and Mesbah, 2021; Sánchez et al., 2020; Zhang et al., 2020; Murphy et al., 2020; Bagherzadeh et al., 2020; Wang et al., 2020; Garcia et al., 2020; Radu and Nadi, 2019; Zhang et al., 2018; Gao et al., 2018; Liu et al., 2019; Ohira et al., 2015; Lin et al., 2015; Le Goues et al., 2015, 2015; Zaman et al., 2012; Jalbert et al., 2011; Ku et al., 2007; Chen et al., 2021b; Selakovic and Pradel, 2017; Gunawi et al., 2014; Ho et al., 2023; de Santana et al., 2024; Azad et al., 2023; Zhao et al., 2023b; Wang et al., 2023c, 2022b; Shen et al., 2021; Wang et al., 2021a; Eghbali and Pradel, 2020; Franco et al., 2017; Wang et al., 2017; Tan et al., 2018) 60
Static Analysis (Senanayake et al., 2023; Liang et al., 2022; Zheng et al., 2021; Kamienski et al., 2021; Ferenc et al., 2020; Tan et al., 2017; Wang et al., 2023b) 7
Machine Learning (Song et al., 2023; Avula et al., 2023; Xu et al., 2019) 3
Others (Ruan et al., 2024; Chen et al., 2023b; Keshavarz and Nagappan, 2022; Bui et al., 2022; Pereira et al., 2022; Vieira et al., 2019; Tomassi et al., 2019; Madeiral et al., 2019; Gkortzis et al., 2018) 9
@@
By far the most common way for dataset creators to categorize the defects in their datasets is through manual examination, with 60 (60/132) datasets falling into this category. Similar to manual validation, manual categorization allows researchers to operate reasonably efficiently and flexibly. Indeed, it is not uncommon for datasets to employ both manual validation and manual categorization, (e.g., ECench (Kim et al., 2022) and ExcePy (Zhang et al., 2022b)), with 43 (43/132) datasets in this group.
Defect categorization need not be entirely automatic or entirely manual. For instance, Ohira et al. (2015) label “surprise” and “dormant” bugs automatically, since the definitions they use are easy to detect, while manually labeling “blocking,” “security,” “performance,” and “breakage” bugs.
@@
Seven datasets in the survey use static analysis to classify defects. For instance, Senanayake et al. (2023) run two static analyzers, Qark (qar, 2025) and MobSF (mob, 2025), on the APKs and source files of Android applications and store the analysis results in their dataset. Categorizing defects based on static analysis has some of the same benefits and drawbacks of static analysis-based collection, namely a fairly high precision, but a limited scope of possible categories and the possibility of false negatives (i.e., bugs with a missing classification).
@@
In total, three of the surveyed datasets classify their entries using machine learning approaches. Song et al. (2023) train a neural network to classify bugs given a bug report, fixing commit message, bug-fix location, and bug-inducing location. Avula et al. (2023) use an off-the-shelf LLM to generate a description of what each bug fix in their dataset does. Xu et al. (2019) use latent Dirichlet allocation (Blei et al., 2003) (a kind of Bayesian network) to extract topic keywords from pull requests and add them to each defect’s metadata.
5. Availability and Usability

### v2

@@ 4.2.1. Defect Collection
There are six datasets in the survey that use static analysis to collect defects. To this end, the dataset creators run existing or newly created static analyses on candidate code. As an example of the former, DBLP:conf/icse/ZhengPLBEYLMS21 create D2A by running the static analyzer Infer on pairs of commits, looking for detected issues that are present in one commit but absent in the next. As an example of the latter, ManySStuBs4J (DBLP:conf/msr/KarampatsisS20) uses a custom static analysis approach that compares the ASTs of candidate buggy and fixed programs and searches for a set of specific bug-fix patterns. An advantage of using static analysis is that, depending on the specific tools or techniques used, one can achieve better precision than strategies such as keyword matching (i.e., fewer non-bugs classified as bugs). The main disadvantage is that relying on static analysis to separate bugs from non-bugs will limit a dataset’s scope to the abilities of the static analysis. Many static analyzers only detect specific classes of bugs (e.g., null pointer bugs), and even more generalist analyzers like Cppcheck will necessarily fail to catch some bugs (Habib and Pradel, 2018).
@@
There are three datasets in the survey that use machine learning techniques as part of their data collection processes. DBLP:conf/icse/WangHGWC024 input candidate patches into LLMs, asking the models to evaluate whether those patches fix software vulnerabilities. Their dataset also illustrates that extra care must be taken to ensure that LLM output is sound, which they achieve by using LLMs in conjunction with static analysis tools and by including only defects where the two agree. Similarly, DBLP:conf/icse/ZhengPLBEYLMS21 and DBLP:journals/jss/WuZCWM20 use machine learning models to distinguish between defects and non-defects based on semantic information, such as bug reports or commit messages.
@@ 4.2.2. Validation
Manual Inspection Full (DBLP:journals/ese/TimperleyHSDW24; DBLP:journals/tosem/LiuMC24; DBLP:journals/tosem/ZhangCWCLMMHL24; DBLP:conf/icse/CuiD0WSZWYXH0024; DBLP:conf/issta/YuXZ0LS24; DBLP:conf/issta/XuG024; DBLP:conf/ease/WaseemDA0M24; DBLP:conf/kbse/DasAM24; DBLP:conf/acl/TianYQCLPWHL0024; DBLP:conf/msr/SilvaSM24; DBLP:conf/icst/LeeKYY24; DBLP:journals/jss/ZhaoMLZ23; DBLP:conf/icst/KimH23; DBLP:journals/ese/MorovatiNKJ23; DBLP:conf/msr/ApplisP23; DBLP:conf/icse/GuanXLLB23; DBLP:conf/issta/XiongX0SWWP0023; DBLP:conf/icse/LiangLSSFD22; DBLP:conf/msr/KimKL22; DBLP:conf/wcre/ZhangYYCYZ22; DBLP:journals/pacmpl/PaltenghiP22; DBLP:conf/issre/LiuZDHDM022; DBLP:conf/msr/WendlandSMMHMRF21; DBLP:journals/stvr/GyimesiVSMBFM21; DBLP:conf/cgo/YuanLLLLX21; DBLP:conf/icse/PaulTB21a; DBLP:conf/issre/Du0MZ21; DBLP:conf/icsm/ZhangXL20; DBLP:conf/hotsos/MurphyBSR20; DBLP:conf/sigsoft/WidyasariSLQPTT20; DBLP:journals/pacmpl/0001FSK20; DBLP:conf/msr/WangBJS20; DBLP:conf/icse/BentonGZ19; DBLP:conf/msr/RaduN19; DBLP:conf/msr/RiganelliMMM19; DBLP:conf/wcre/DelfimUMM19; DBLP:conf/msr/PontaPSBD19; DBLP:conf/dsa/XuYWA19; DBLP:conf/msr/SahaLLYP18; DBLP:conf/issta/ZhangCCXZ18; DBLP:journals/ese/LiuWWXCWYZ19; DBLP:conf/kbse/LinMZCZ15; DBLP:conf/issta/JustJE14; DBLP:conf/kbse/KuHCL07; DBLP:conf/icse/ChenYLCLWL21; DBLP:conf/cloud/GunawiHLPDAELLM14; DBLP:conf/icsm/HoMI0SKNR23; DBLP:conf/uss/XuLDDLWPM23; DBLP:journals/tosem/SantanaNAA24; DBLP:conf/msr/AzadIHR23; DBLP:conf/qsw/ZhaoWLLZ23; DBLP:conf/kbse/WangWCCY22; DBLP:conf/sigsoft/ShenM0TCC21; DBLP:conf/kbse/EghbaliP20; DBLP:conf/kbse/FrancoGR17; DBLP:conf/kbse/WangDGGQYW17; DBLP:conf/icse/TanDGR18; DBLP:conf/icse/HumbatovaJBR0T20; DBLP:conf/sigsoft/GaoDQGW0HZW18; DBLP:conf/msr/BeyerGKRR24; DBLP:journals/ese/TambonNAKA24; DBLP:journals/tse/ZhaoXBCL23) 62
Partial (DBLP:conf/icse/WangHGWC024; DBLP:conf/kbse/ZhongLZGWGL23; DBLP:conf/kbse/AvulaVM23; DBLP:conf/issta/Song0NW0DM22; DBLP:conf/apsec/AkimovaBDKKMM21; DBLP:conf/msr/KimKL21; DBLP:conf/icse/ZhengPLBEYLMS21; DBLP:conf/icse/Makhshari021; DBLP:journals/jss/FerencGGTG20; DBLP:journals/jss/WuZCWM20; DBLP:journals/ese/DAmbrosLR12; DBLP:conf/msr/ZamanAH12; DBLP:journals/infsof/WangBWYGS23; DBLP:journals/jss/WangCHZBZ23; DBLP:conf/compsac/GuWL0019; DBLP:conf/kbse/RomanoLK021; DBLP:conf/msr/OliverDAH24; DBLP:journals/pacmpl/DrososSAM024; DBLP:journals/tosem/TufanoWBPWP19) 19
Defect Reproduction (DBLP:journals/ese/TimperleyHSDW24; DBLP:conf/issta/YuXZ0LS24; DBLP:conf/acl/TianYQCLPWHL0024; DBLP:conf/iclr/JimenezYWYPPN24; DBLP:conf/msr/PramodSTSW24; DBLP:conf/msr/SilvaSM24; DBLP:conf/icst/LeeKYY24; DBLP:journals/jss/ZhaoMLZ23; DBLP:conf/icst/KimH23; DBLP:conf/icse/SaavedraSM24; DBLP:journals/compsec/BrustSG23; DBLP:journals/tse/JiangLLZCNZHBZ23; DBLP:conf/msr/ApplisP23; DBLP:conf/icse/BhuiyanPVPS23; DBLP:conf/issta/XiongX0SWWP0023; DBLP:conf/wcre/ZhangYYCYZ22; DBLP:conf/msr/BuiSF22; DBLP:conf/issta/Song0NW0DM22; DBLP:conf/msr/WendlandSMMHMRF21; DBLP:journals/stvr/GyimesiVSMBFM21; DBLP:conf/cgo/YuanLLLLX21; DBLP:conf/sigsoft/WidyasariSLQPTT20; DBLP:conf/icse/BentonGZ19; DBLP:conf/icse/DmeiriTWBLDVR19; DBLP:conf/msr/RiganelliMMM19; DBLP:conf/wcre/DelfimUMM19; DBLP:conf/msr/SahaLLYP18; DBLP:conf/issta/ZhangCCXZ18; DBLP:conf/msr/GaoYJLYZ18; DBLP:journals/tse/GouesHSBDFW15; DBLP:conf/issta/JustJE14; DBLP:conf/hotpar/JalbertPPS11; DBLP:conf/se/SelakovicP17; DBLP:conf/uss/XuLDDLWPM23; DBLP:conf/issta/WuJPLD0BS23; DBLP:conf/wcre/ReyesGSBM24; DBLP:journals/access/AntalVKMHF24; DBLP:journals/ese/TambonNAKA24; DBLP:conf/kbse/XuZL23) 39
@@
Dataset validation is about ensuring that the defects included in a dataset are actual defects. This is an important step for making a useful dataset, since a high false positive rate (i.e., many non-bugs classified as bugs) can taint further research that uses a dataset. Even so, not all dataset creators take this step, and there are legitimate reasons for doing so. For instance, a dataset might be so large that validation becomes impractical (e.g., (DBLP:conf/edcc/AndradeLV24; DBLP:conf/msr/CsuvikV22; DBLP:conf/msr/MahbubSR23; DBLP:conf/raid/0001DACW23; DBLP:conf/aina/PonteRM23a)), or a dataset’s data collection approach might already have a low false positive rate (e.g., (DBLP:conf/msr/LiuHLZCSHM24; DBLP:journals/jss/CaiLZZS24; DBLP:conf/kbse/SongWCCLLWP23; DBLP:journals/tse/ChenXLGLC22; DBLP:conf/icse/TanYYMR17; DBLP:conf/sp/Dolan-GavittHKL16)). Table 8 shows common validation methodologies and datasets using them.
@@
Some dataset creators, such as DBLP:conf/icst/LeeKYY24, DBLP:conf/msr/WangBJS20, and DBLP:conf/msr/SilvaSM24, opt to manually inspect the code and description of every bug in their datasets (i.e., full manual inspection). This allows researchers to construct a dataset with a low false positive rate and to make qualitative judgments about what bugs should be included, without spending the time and resources needed for manual defect collection. Full manual inspection is quite popular among the datasets surveyed, with 62 (62/151) datasets employing it. The largest (DBLP:conf/acl/TianYQCLPWHL0024) and smallest (DBLP:conf/icst/KimH23) datasets that employ this strategy have 4,253 and 18 entries respectively, with a median of 260 entries.
Conversely, some dataset creators choose to only manually inspect a subset of the defects in their dataset (i.e., partial manual inspection). This is less common than full inspection, with only 19 (19/151) datasets employing it. However, it is attractive for particularly large datasets where full inspection is infeasible, such as D2A (DBLP:conf/icse/ZhengPLBEYLMS21), the largest dataset that uses this approach, which has over 1.3 million entries. The smallest such dataset (DBLP:conf/msr/ZamanAH12) still contains 191 entries, with the median size being 5,565 entries.
Typically, partial inspection has no effect on the actual construction of the dataset, but is performed to demonstrate the accuracy of an automated methodology (DBLP:conf/icse/ZhengPLBEYLMS21). However, there is at least one case where partial inspection is used to filter out defects from a dataset. DBLP:conf/kbse/ZhongLZGWGL23 automatically classify each bug in their dataset into one of multiple fault categories. They then manually analyze 50 bugs from each category to ascertain that category’s accuracy, and only include bugs from categories with over 90% accuracy in the final dataset.
@@
Among our surveyed datasets, 39 (39/151) involve defect reproduction, and only one dataset attempts to reproduce only a subset of its defects, while the rest aim for full reproduction. This process is often facilitated by sources that provide a rich environment and execution context, such as continuous integration (CI) systems. For example, BugSwarm (DBLP:conf/icse/DmeiriTWBLDVR19) leverages CI configurations and historical build logs to automatically reproduce bugs in Docker-based environments, thus making the automation of reproduction feasible.
However, automation has limitations. Many defects are not covered by existing tests, or their reproduction may require complex setup steps beyond test execution (e.g., user interface-related bugs in AndroR2 (DBLP:conf/msr/WendlandSMMHMRF21)). Still, automated reproduction enables the construction of much larger datasets than manual validation strategies. The largest dataset in this category (DBLP:journals/compsec/BrustSG23) contains approximately 168,000 entries. The median size, 225 entries, is comparable to that of datasets which use full manual inspection.
@@ 4.2.3. Categorization
Manual Categorization (DBLP:journals/ese/TimperleyHSDW24; DBLP:journals/tosem/LiuMC24; DBLP:journals/tosem/ZhangCWCLMMHL24; DBLP:journals/tosem/JiangJWMLZ24; DBLP:conf/icse/CuiD0WSZWYXH0024; DBLP:conf/issta/YuXZ0LS24; DBLP:conf/issta/XuG024; DBLP:conf/ease/WaseemDA0M24; DBLP:conf/kbse/DasAM24; DBLP:conf/qrs/YuW24; DBLP:conf/kbse/SongWCCLLWP23; DBLP:journals/jss/ZhaoMLZ23; DBLP:conf/icst/KimH23; DBLP:journals/ese/MorovatiNKJ23; DBLP:conf/msr/ApplisP23; DBLP:conf/icse/GuanXLLB23; DBLP:conf/issta/XiongX0SWWP0023; DBLP:conf/icse/LiangLSSFD22; DBLP:conf/msr/KimKL22; DBLP:conf/wcre/ZhangYYCYZ22; DBLP:journals/tse/ChenXLGLC22; DBLP:journals/pacmpl/PaltenghiP22; DBLP:conf/issre/LiuZDHDM022; DBLP:conf/cgo/YuanLLLLX21; DBLP:conf/icse/PaulTB21a; DBLP:conf/msr/KamienskiPBH21; DBLP:conf/issre/Du0MZ21; DBLP:conf/icse/Makhshari021; DBLP:journals/access/SanchezDMS20; DBLP:conf/icsm/ZhangXL20; DBLP:conf/hotsos/MurphyBSR20; DBLP:journals/pacmpl/0001FSK20; DBLP:conf/msr/WangBJS20; DBLP:conf/icse/GarciaF0AXC20; DBLP:conf/msr/RaduN19; DBLP:conf/issta/ZhangCCXZ18; DBLP:conf/msr/GaoYJLYZ18; DBLP:journals/ese/LiuWWXCWYZ19; DBLP:conf/msr/OhiraKYYMLFHIM15; DBLP:conf/kbse/LinMZCZ15; DBLP:journals/tse/GouesHSBDFW15; DBLP:journals/tse/GouesHSBDFW15; DBLP:conf/msr/ZamanAH12; DBLP:conf/hotpar/JalbertPPS11; DBLP:conf/kbse/KuHCL07; DBLP:conf/icse/ChenYLCLWL21; DBLP:conf/se/SelakovicP17; DBLP:conf/cloud/GunawiHLPDAELLM14; DBLP:conf/icsm/HoMI0SKNR23; DBLP:journals/tosem/SantanaNAA24; DBLP:conf/msr/AzadIHR23; DBLP:conf/qsw/ZhaoWLLZ23; DBLP:conf/wcre/WangZRLJ23; DBLP:conf/kbse/WangWCCY22; DBLP:conf/sigsoft/ShenM0TCC21; DBLP:conf/sigsoft/WangLX0S21; DBLP:conf/kbse/EghbaliP20; DBLP:conf/kbse/FrancoGR17; DBLP:conf/kbse/WangDGGQYW17; DBLP:conf/icse/TanDGR18; DBLP:conf/icse/HumbatovaJBR0T20; DBLP:conf/kbse/RomanoLK021; DBLP:conf/msr/AmannNNNM16; DBLP:conf/sigsoft/GaoDQGW0HZW18; DBLP:conf/msr/BeyerGKRR24; DBLP:journals/ese/TambonNAKA24; DBLP:journals/pacmpl/DrososSAM024; DBLP:journals/tse/ZhaoXBCL23) 68
Static Analysis (DBLP:conf/secrypt/SenanayakeKAP023; DBLP:conf/icse/LiangLSSFD22; DBLP:conf/icse/ZhengPLBEYLMS21; DBLP:conf/msr/KamienskiPBH21; DBLP:journals/jss/FerencGGTG20; DBLP:conf/icse/TanYYMR17; DBLP:journals/jss/WangCHZBZ23; DBLP:conf/fie/SandersWA24) 8
Machine Learning (DBLP:conf/kbse/SongWCCLLWP23; DBLP:conf/kbse/AvulaVM23; DBLP:conf/dsa/XuYWA19) 3
Others (DBLP:conf/kbse/RuanLZL24; DBLP:conf/raid/0001DACW23; DBLP:conf/msr/KeshavarzN22; DBLP:conf/msr/BuiSF22; DBLP:conf/prdc/PereiraAV22; DBLP:conf/promise/VieiraSRG19; DBLP:conf/icse/DmeiriTWBLDVR19; DBLP:conf/wcre/DelfimUMM19; DBLP:conf/msr/GkortzisMS18; DBLP:conf/sigsoft/Zhang0C0Z19; DBLP:conf/wcre/ReyesGSBM24) 11
@@
By far the most common way for dataset creators to categorize the defects in their datasets is through manual examination, with 68 (68/151) datasets falling into this category. Similar to manual validation, manual categorization allows researchers to operate reasonably efficiently and flexibly. Indeed, it is not uncommon for datasets to employ both manual validation and manual categorization, (e.g., ECench (DBLP:conf/msr/KimKL22) and ExcePy (DBLP:conf/wcre/ZhangYYCYZ22)), with 48 (48/151) datasets in this group.
Defect categorization need not be entirely automatic or entirely manual. For instance, DBLP:conf/msr/OhiraKYYMLFHIM15 label “surprise” and “dormant” bugs automatically, since the definitions they use are easy to detect, while manually labeling “blocking,” “security,” “performance,” and “breakage” bugs.
@@
Eight datasets in the survey use static analysis to classify defects. For instance, DBLP:conf/secrypt/SenanayakeKAP023 run two static analyzers, Qark and MobSF , on the APKs and source files of Android applications and store the analysis results in their dataset. Categorizing defects based on static analysis has some of the same benefits and drawbacks of static analysis-based collection, namely a fairly high precision, but a limited scope of possible categories and the possibility of false negatives (i.e., bugs with a missing classification).
@@
In total, three of the surveyed datasets classify their entries using machine learning approaches. DBLP:conf/kbse/SongWCCLLWP23 train a neural network to classify bugs given a bug report, fixing commit message, bug-fix location, and bug-inducing location. DBLP:conf/kbse/AvulaVM23 use an off-the-shelf LLM to generate a description of what each bug fix in their dataset does. DBLP:conf/dsa/XuYWA19 use latent Dirichlet allocation (Blei et al., 2003) (a kind of Bayesian network) to extract topic keywords from pull requests and add them to each defect’s metadata.
RQ2 Summary The majority of software defect datasets are sourced from software development processes, with issue reports and version control systems being most common. A smaller number rely on pre-existing datasets, vulnerability databases, developer Q&A forums, coding assignments, and CI/CD systems. Some datasets employ defect synthesis techniques to generate synthetic defects for controlled experimentation. Dataset construction generally involves three stages: defect collection, validation, and categorization. Collection is most often performed through automated mining strategies such as keyword matching (64/151) or mapping (42/151) between issue reports and commits. Fewer datasets rely on manual collection, external oracles, the SZZ algorithm, static analysis, or machine learning techniques. Validation and categorization practices vary widely, though manual inspection remains most common in both stages.
5. Availability and Usability (RQ3)